You roll two fair dice, one green and one red.

    (a) Are the outcomes on the dice independent?
Yes No      

    (b) Find P(5 on green die and 3 on red die). (Enter your answer as a fraction.)
2

    (c) Find P(3 on green die and 5 on red die). (Enter your answer as a fraction.)
3

    (d) Find P((5 on green die and 3 on red die) or (3 on green die and 5 on red die)). (Enter your answer as a fraction.)
4
Answer #1

(a) yes


(b) 1/6 * 1/6 = 1/36


(c) 1/6 * 1/6 = 1/36


(d) That's (b) + (c) = 1/18
Answer #2

a) yes

b) (1/6)^2 = 1/36


c) 1/36

d)2/36 = 1/18